OpenStudy (anonymous):

3(2x+7)=45 A 19 B 64 C 4 D 11 -23= -2a - 10 +a a a=31 b a= 13 c a=-11 d a= 5

OpenStudy (anonymous):

Please Help me

hartnn (hartnn):

3(2x+7)=45 start by dividing 3 on both sides, what u get ?

OpenStudy (anonymous):

so divide 3 an 7 an then 3 an 45

OpenStudy (anonymous):

6x + 21 = 45 6x = 24 x = 4 -23 = -2a - 10 -1a -13 = -a 13 = a

OpenStudy (anonymous):

confused like i know how to do it but every time i do i get some thing wrong an cant find where i went wrong i did it exactly like your DemolisionWolf

hartnn (hartnn):

3(2x+7)=45 dividing both sides by 3 3(2x+7)/3 =45/3 the 3's on left get cancelled, so, 2x+7 = 15 got this ?

OpenStudy (anonymous):

yea

hartnn (hartnn):

2x+7 =15 subtract 7 from both sides, what u get ?

hartnn (hartnn):

2x= 15-7 2x=8 so x=...?